OpenStudy (anonymous):

Help!! The roots of a function are at x=-1, x=1/4i, and x=-1/4i. Which statement is true about this function? A. No such function exists B. The graph of the function crosses the x-axis only one C. The graph of the function is increasing over any interval. D. The expanded form of the function will have an imaginary coefficient.

OpenStudy (zehanz):

Well, A is wrong, because you can define:\[f(x)=(x+1)(x-\frac{ 1 }{ 4 }i)(x+\frac{ 1 }{ 4 }i)\]so such a function does exist. What do you think, B, C or D as answer?

OpenStudy (anonymous):

D?

OpenStudy (zehanz):

How many times do you think the graph crosses the x-axis?

OpenStudy (anonymous):

Once..?

OpenStudy (zehanz):

It must, because -1 is a zero. It also must be the only time, otherwise there would be more real zeros! B is the right answer. BTW if you multiply out the formula of f, you get:\[f(x)=(x+1)(x^2 -\frac{ 1 }{ 16 }i^2)=(x+1)(x^2+\frac{ 1 }{ 16 }) = ...\]See? There are no more real zeros, also there is no imaginary coefficient.

OpenStudy (anonymous):

Thank you so much! Sorry for being such a hassle!!

OpenStudy (zehanz):

No problem!